Hypervelocity bullet used in 0.22 long rifle usually weigh around 2.1 g and can have a speed of 550 m/s. what must be the speed of a 75 kg man to match the momentum of this bullet?


1
Expert's answer
2021-04-12T06:54:43-0400

"m_1v_1=m_2v_2,"

"v_2=v_1\frac{m_1}{m_2},"

"v_2=550\cdot \frac{0.0021}{75}=15.4~\frac{mm}{s}."